OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)) helppppppp

OpenStudy (gibbs):

6 x 5 x 3 = 90

OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)). -40x+24 -40x+31 -40x+64 -40x+70

OpenStudy (mimi_x3):

Ok so our g(x) replaces every x in f(x)=-4x+7 f(g(x))=-4(10x-6)+7 Distribute and then combine like terms

OpenStudy (mimi_x3):

Distribute the -4 meaning mutiply 10x and -6 by -4